Is there a path from start to finish?

You are given a graph where two nodes are labeled “start” and “finish.” The problem is to determine whether there is a path between these two nodes. But the graph is layered, and you can only see one layer at a time!

2 3 4 5 6 7 8 9 10 11 12 13 14